Question : evaluate each of the following arithmetic series
1, -3, -7, ..., -75

Answer by wangsol(2)

1, -3, -7, ..., -75
a1= 1
a2= -3
d=-3-1=-4
an=-75
n=?
using the formula:
an=a1+(n-1)*d
-75=1+(n-1)*-4
-75=1-4n+4
-75=5-4n
-80=-4n
n=(-80)/(-4)
n=20 terms
1,-3,-7,-11,-15,-19,-23,-27,-31,-35,-39,-43,-47,-51,-55,-59,-63,-67,-71,-75
d=-4
1+(-4)=-3
(-3)+(-4)=-7
(-7)+(-4)=-11
(-11)+(-4)=-15
and so on...
(-71)+(-4)=-75